OpenStudy (anonymous):

How do i solve 5x+2y=1 2x-3y=8

OpenStudy (anonymous):

using the elimination method

OpenStudy (anonymous):

U can solve by either subsitution or elimnating one of the term:: solve for x, and plug into other equation: or multiply a number by one of the equation that will eliminate 1 term when added.

OpenStudy (anonymous):

can you show me? please

OpenStudy (anonymous):

so elimination: let solve the 2nd one, 2x-3y=8, solve for x, u get x=4+(3/2)y, plug back to first equation: 5(4+(3/2)y)+2y=1, solve for y, u get y=-2, and plug that back in 2nd equation: and u get (2x-3(-2))=8, so x=1

OpenStudy (anonymous):

im having trouble showing my work

OpenStudy (anonymous):

i understand now but not sure that i'm writing it correctly

OpenStudy (anonymous):

i dont know what im forgetin to do, my work just doest look wright. it looks like i forgot somethin

OpenStudy (anonymous):

mm?

OpenStudy (anonymous):